whts the value of cos 4x can you reply it back to the mail id deepan.nj at gmail.com

Grade:12th pass

2 Answers

Amit
5 years ago
Hi there.. I didn`t got what you are trying to ask. Whether you formula or something else.. For formula just exapnd it as 2cos²2x -1 and again use formula of cos2x.
Md hasan
5 years ago
If we do the differentiate with respect to x...Answer will be -4sinx....Dy/dx=d(cos4x)/dx=(-sinx)×4=-4sinx